OpenStudy (anonymous):

(x-2)^2 -5 =11 Solve quadratic function with square roots

OpenStudy (alexandervonhumboldt2):

first subtract 11 from both sides: we get (x-2)^2-16 16=4^2 use difference of squares: a^2-b^2=(a+b)*(a-b) a=x-2 b=4 substitute: (x+2)*(x-6)=0 equal both ( ) to 0: x+2=0 x-6=0 solve each equation and it will be your roots